HUBBLE SPACE TELESCOPE DAILY REPORT #1804 PERIOD COVERED: 0000Z (UTC) 01/28/97 - 0000Z (UTC) 01/29/97 Daily Status Report as of 029/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6419 (Testing Unified Models With a Complete = Sample of Seyfert Galaxies) The WF/PC-2 was used to make observations of the Seyfert galaxy = MCG-6-30-1 to help build a complete survey of the Seyfert class of = galaxies. This will help test the current model for Seyfert class = galaxies. This model proposes that all Seyferts have Seyfert-1 nucleus, = which is surrounded by a torus of gas and dust. The observations were = completed as planned and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: = Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 1.3 Completed GHRS/1/2 6617 (The ISM Toward Nearby High Velocity = Stars) The High Resolution Spectrograph (Both Sides One and Two ) was = used to make spectrographic observations of the high velocity star HD26965= in order to study the inter-stellar medium (ISM) in the direction toward = this star. This is to determine the accurate deuterium to hydrogen ratio = in this local interstellar gas. A wavelength calibration image was also = taken. The observations were executed as scheduled, and no anomalies = were noted. 1.4 Completed WF/PC-2 7553 (Pure Parallel Archive Survey II) WF/PC-2 parallel observations were obtained in the ultraviolet in = search for interesting objects. This program is designed to find = extragalactic sources with near ultraviolet excesses (3000 A0) outside = the galactic plane and far ultraviolet sources (1600 A0) in the galactic = plane. The procedure was executed as scheduled, and no anomalies were = reported. 1.5 Completed FOS/BL/GHRS/1 5947 (The Relationship Between Galaxies = and Lyman-Alpha Absorbers at Low Redshift) The Faint Object Spectrograph (Blue Side) and the High Resolution = Spectrograph (Side One) were used to make observations of Quasar = Q1214+180 to study the relation ship between its galaxy and Lyman-Alpha = absorbers at low redshift. The relationship between the galaxy and low = column density, metal poor quasar absorbers is poorly understood. A = wavelength calibration image was also taken. A Loss-of-Lock occurred FGS = #1 and #3 at 028/1413Z for about two minutes then re-acquired lock. HSTAR = #5887 was written. The observations were executed as scheduled, and no = anomalies were noted. 1.6 Completed FOS/RD 6910 (FOS Cycle 6: Focus, X-Pitch, and Y-Pitch) The Faint Object Spectrograph (Red Side) was used to make focus = calibration observations of the X and Y pitch values to test the system's = focus. This will allow for optimal focus determination of the = spectrograph. A wavelength calibration image was also taken. The = observations were executed as scheduled, and no anomalies were noted. 1.7 Completed FOS/RD 6801 (C and N Production and Pollution Mechanism = in Low-Metallicity) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of the dwarf galaxy NGC-4861 to measure its = N, O and C emission lines in its H II regions. The observations were = completed as planned and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 9 Successful : 9 Loss-of-lock occured at 028/1413Z, HSTAR #5887 was written. See Note 1.5 in this report. 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: JIS #2 was completed last night as planned. The JIS #3 Debriefing = is scheduled this morning at 9am EST on the Prime Ops Loop. The Real-time HST flight operations remains at SAMS until the = completion of JIS #4, which is scheduled for completion at 5:30pm EST = today. 3.0 SIGNIFICANT FORTHCOMING EVENTS: JIS #4 (Short HST Deploy) is scheduled to start this morning at = 11:30am and to last for about six hours. /DMH